# Nearly Optimal Bayesian Inference for Structural Missingness

### v1

@@ Abstract
Structural missingness breaks ’just impute and train’: values can be undefined by causal or logical constraints, and the mask may depend on observed variables, unobserved variables (MNAR), and other missingness indicators. It simultaneously brings (i) a catch-22 situation with causal loop, prediction needs the missing features, yet inferring them depends on the missingness mechanism, (ii) under MNAR, the unseen are different, the missing part can come from a shifted distribution, and (iii) plug-in imputation, a single fill-in can lock in uncertainty and yield overconfident, biased decisions. In the Bayesian view, prediction via the posterior predictive distribution integrates over the full model posterior uncertainty, rather than relying on a single point estimate. This framework decouples (i) learning an in-model missing-value posterior from (ii) label prediction by optimizing the predictive posterior distribution, enabling posterior integration. This decoupling yields an in-model almost-free-lunch: once the posterior is learned, prediction is plug-and-play while preserving uncertainty propagation. It achieves SOTA on 43 classification and 15 imputation benchmarks, with finite-sample near Bayes-optimality guarantees under our SCM prior.
@@ 3 Bayesian Inference with Structural Missingness
Note that $X_{m}^{c}$ explicitly includes the mask $M$, ensuring that the ”observation” includes information about which variables are observed. $X_{m}$ and $X_{m}^{c}$ are still well-defined random variables to perform Bayesian inference. The formal proof that $X_{m}$ and $X_{m}^{c}$ constitute valid random variables on the appropriate disjoint union spaces is provided in Appendix C.
@@ 6.1 Risk Minimization and Approximation
This is proved in the Appendix Sec. D.
@@ 6.2 In-Model Reduction of Plug-in Imputation and MNAR Bias
This is proved in the Appendix Sec. E.
@@
This is proved in the Appendix Sec. F.
@@ 6.3 Sample Complexity Advantage of Decoupled Inference
Theorem 6.7 gives a sample-complexity advantage (ICL difficulty proxy) for learning $G$ (missing-feature posterior) and $H$ (conditional predictor) separately rather than end-to-end in finite-context / ICL. Appendix Sec. H provides the supporting error decompositions and proof chain (Theorem H.5; Corollary H.6).
@@
Assume the conditional predictor regularity in Appendix Sec. H so that the posterior-mismatch term is controlled by $L_{h}\,W_{1}(\cdot,\cdot)$ (Lemma H.3). Then Corollary H.6 implies that to reach total error $\epsilon$ it suffices to learn $G$ to Wasserstein accuracy $\epsilon/(2L_{h})$ and $H$ to conditional-prediction error $\epsilon/2$, capturing the required error propagation.
@@
Thus, when $L_{g}$ is large and context is limited (large $L_{\mathrm{cpl}}$), learning $G$ and $H$ separately can be more sample-efficient than learning the coupled end-to-end map; see Appendix Sec. H.
@@ 7 Experiments
XGBoost + MOT-MLP $0.8073\pm 0.1216$ 17.50 $132.754\pm 0.776$
PFN-NSM $0.8065\pm 0.1219$ 16.00 $0.061\pm 0.003$
XGBoost + IGRM $0.8048\pm 0.1127$ 17.80 $11.809\pm 0.156$
@@ Appendix A Related Work
Structural Missingness. While classic MCAR/MAR/ MNAR definitions (Rubin, 1976) cover probabilistic missingness, structural missingness deals with logically undefined values. Recent works have modeled this using graph-based approaches or separate categories, but often lack a unified Bayesian treatment. (Mohan et al., 2013) proposes and proved that dealing with missing data is a causal inference problem in a unified identification framework. However, due to lack of integration of methods and tools, models with simplified assumptions without considering selection bias caused by not explicitly differentiating $P(x|M=0)$ and $P(x|M=1)$ still dominate the field with plausible performance (Muzellec et al., 2020; Vo et al., 2024; Zhang et al., 2025). Though the structure is not identifiable in general without extra assumptions, PFNs can offer an “almost-free-lunch” in terms of efficient amortized inference within an assumed SCM prior: when the prior renders the relevant posteriors identifiable in-model (Remark 6.6), PFNs can approximate these posteriors and enable practical inference, rather than resolving fundamental MNAR non-identifiability without assumptions.
@@
Prior-Fitted Networks. PFNs (Müller et al., 2022, 2025) and TabPFN (Hollmann et al., 2022) demonstrate the power of learning Bayesian inference on tabular data. These methods have shown state-of-the-art performance on many small or medium-sized datasets since TabPFN-v2 (Hollmann et al., 2025). As concerns about scalability to larger datasets are raised, TabICL (Qu et al., 2025) proposes a distribution-aware feature embedding method to improve scalability, and BETA (Liu and Ye, 2025) adapts a lightweight encoder to align TabPFN to large-scale high-dimensional tabular data. As shown in the technical report for TabPFN-2.5 (Grinsztajn et al., 2025), it achieves dominant performance (87% win rate) on tabular datasets up to 100K samples and 2K features. PFN-based methods also have differentiable properties that can be leveraged to produce feature-importance measures, partial dependence (Rundel et al., 2024), and high-dimensional Bayesian optimization (Müller et al., 2023), and show promising results in settings requiring causal mechanisms, such as causal effect estimation (Balazadeh et al., 2025) and causal discovery (Robertson et al., 2025).
Generative Imputation. In tabular imputation, we aim to sample from a conditional distribution of missing entries given observed entries and a missingness mask, under heterogeneous feature types (continuous, discrete, categorical) and strong cross-feature constraints; in many practical settings this conditional mapping is close to deterministic (e.g., structurally constrained columns and rule-like relations).
@@ B.1 Missing Data Classification Experiments
Datasets (summary).
We summarize the classification datasets below, which are selected from OpenML and grouped by empirical missing rate, including dataset size (NN), inferred feature dimension ($D$), and missingness statistics (total missing entries and overall missing rate). See Tables 4 and 5.
@@ B.2 Missing Data Imputation Experiments
XGBoost + MOT-MLP $0.8073\pm 0.1216$ 17.50 $132.754\pm 0.776$
PFN-NSM $0.8065\pm 0.1219$ 16.00 $0.061\pm 0.003$
XGBoost + IGRM $0.8048\pm 0.1127$ 17.80 $11.809\pm 0.156$
@@
TabPFN + MOT-LIN $0.8011\pm 0.1290$ 23.20 $127.294\pm 3.778$
TabPFN + MOT-MLP $0.8006\pm 0.1311$ 23.40 $128.127\pm 0.776$
TabPFN + TDM $0.8006\pm 0.1294$ 24.70 $123.860\pm 8.327$
CatBoost + Remasker $0.8005\pm 0.1298$ 27.70 $551.288\pm 9.349$
@@
TabPFN + MissForest $0.8002\pm 0.1272$ 23.10 $26.122\pm 1.681$
TabPFN + HyperImpute $0.8001\pm 0.1285$ 23.20 $170.758\pm 0.510$
@@
XGBoost + MissForest $0.7992\pm 0.1190$ 22.00 $30.749\pm 1.682$
CatBoost + IGRM $0.7986\pm 0.1155$ 25.30 $44.152\pm 0.241$
TabPFN (Raw) $0.7985\pm 0.1269$ 23.30 $0.061\pm 0.003$
TabPFN + MCFlow $0.7983\pm 0.1214$ 25.70 $28.563\pm 0.583$
XGBoost + HyperImpute $0.7982\pm 0.1281$ 23.40 $175.385\pm 0.510$
TabPFN + OTM-RR $0.7972\pm 0.1272$ 28.30 $5.241\pm 0.101$
@@
XGBoost + MCFlow $0.7935\pm 0.1187$ 21.30 $33.190\pm 0.584$
CatBoost + MCFlow $0.7901\pm 0.1161$ 25.10 $65.533\pm 0.612$
@@
Appendix C The Random Variables are Well-Defined
@@
Appendix E Assumptions and Proof of Theorem 6.3
Assumption E.1 (Posterior approximation and predictor regularity).
@@
We provide a fully detailed proof of Theorem 6.3. Throughout this section, we fix $(X_{m}^{c},D_{m}^{c})$ and use the shorthand $\mu:=P^{\star}(\cdot\mid X_{m}^{c},D_{m}^{c})$ and $\nu:=Q_{\theta}(\cdot\mid X_{m}^{c},D_{m}^{c})$ introduced in Assumption E.1.
Lemma E.2 (Tower property representation).
@@
Lemma E.3 (Error decomposition).
@@
By Lemma E.2, $T_{g}^{\star}=\mathbb{E}_{\mu}[h^{\star}(X_{m})]$. Therefore,
@@
Lemma E.4 (Lipschitz test functions and Wasserstein-1).
@@
Assumption E.1(A0) ensures $W_{1}(\mu,\nu)$ is well-defined. By Kantorovich–Rubinstein duality,
@@
The second term is bounded by Assumption E.1(A3), yielding $\left|\mathbb{E}_{\nu}[h^{\star}-\hat{h}_{\phi}]\right|\leq\varepsilon_{\text{pred}}$. For the first term, Lemma E.4 gives $\left|\mathbb{E}_{\mu}[h^{\star}]-\mathbb{E}_{\nu}[h^{\star}]\right|\leq L\,W_{1}(\mu,\nu)$, and Assumption E.1(A2) implies $W_{1}(\mu,\nu)\leq\varepsilon_{\text{post}}$. Combining yields
@@
Appendix H Posterior-Predictive Approximation via Decoupled Heads
@@
Lemma H.1 (Posterior-predictive marginalization).
@@
Lemma H.2 (Decoupled approximation error decomposition).
@@
Lemma H.3 (Bounding posterior mismatch by $W_{1}$).
Assume that for fixed $(x_{m}^{c},D_{m}^{c})$ and for any $\varphi$ with $\|\varphi\|_{\infty}\leq 1$, the map $x_{m}\mapsto\mathbb{E}_{y\sim H(\cdot\mid x_{m},x_{m}^{c},D_{m}^{c})}[\varphi(y)]$ is $L_{h}$-Lipschitz w.r.t. $\|\cdot\|_{2}$. Then the first term in Lemma H.2 is bounded by
@@
Lemma H.4 (From conditional KL to bounded test functions (Pinsker)).
@@
Proposition H.5 (Why decoupling helps approximate Bayes posterior predictive with limited context).
Under the identifiability condition in Remark 6.6 (so that $P^{\star}(x_{m}\mid x_{m}^{c},D_{m}^{c})$ is well-defined within the model class), Lemmas H.2H.3 show that approximating the Bayes posterior predictive reduces to separately learning (i) the posterior over missing features ($G$) and (ii) the conditional predictor ($H$). Moreover, Lemma H.4 shows how PFN-style training that minimizes conditional cross-entropy (equivalently, conditional KL in the realizable case) controls bounded test-function error for the marginalized posterior predictive without explicitly computing the integral in Lemma H.1. Combined with Theorem 6.7, this supports the claim that, with limited context (test-time training / in-context learning), a decoupled multi-task design can be more sample-efficient and thereby more favorable for approximating the Bayes-optimal posterior predictive than directly learning the coupled end-to-end map.
@@
Under Remark 6.6, the Bayes posterior predictive is well-defined in-model. Lemmas H.2H.3 decompose the approximation error into (i) a posterior-mismatch term (bounded by a $W_{1}$ distance for $G$) and (ii) a conditional-prediction term for $H$. Hence, approximating the posterior predictive reduces to learning $G$ and $H$ to control these two terms. Lemma H.4 provides a direct route for the PFN head: small conditional KL to the true posterior predictive implies small bounded test-function error for the marginalized predictor. Theorem 6.7 adds a complementary complexity argument: when the structural component is highly nonlinear (large effective $L_{g}$) and context is limited (large coupling difficulty), learning the decoupled components can be more sample-efficient than learning the coupled end-to-end map in the ICL regime. ∎
Corollary H.6 (Explicit link to complexity via a two-term target).
Under the assumptions of Lemmas H.2H.3, for any $\varphi$ with $\|\varphi\|_{\infty}\leq 1$,
@@
Apply Lemma H.2 and then upper bound the first (posterior approximation) term by Lemma H.3. The remaining term is exactly $\epsilon_{H}(x_{m}^{c},D_{m}^{c})$ by definition. ∎
Appendix I Proof of Theorem 6.7
@@
By Corollary H.6, to ensure total test-function error at most $\epsilon$ it suffices to satisfy
@@
margins, page numbering, etc.) should be kept the same as the main body.

### v2

@@ Abstract
Structural missingness breaks ’just impute and train’: values can be undefined by causal or logical constraints, and the mask may depend on observed variables, unobserved variables (MNAR), and other missingness indicators. It simultaneously brings (i) a catch-22 situation with causal loop, prediction needs the missing features, yet inferring them depends on the missingness mechanism, (ii) under MNAR, the unseen are different, the missing part can come from a shifted distribution, and (iii) plug-in imputation, a single fill-in can lock in uncertainty and yield overconfident, biased decisions. In the Bayesian view, prediction via the posterior predictive distribution integrates over the full model posterior uncertainty, rather than relying on a single point estimate. This framework decouples (i) learning an in-model missing-value posterior from (ii) label prediction by optimizing the predictive posterior distribution, enabling posterior integration. This decoupling yields an in-model almost-free-lunch: once the posterior is learned, prediction is plug-and-play while preserving uncertainty propagation. It achieves SOTA on 43 classification with missing data and 15 imputation benchmarks, with finite-sample near Bayes-optimality guarantees under our SCM prior.
@@ 1 Introduction
Code and scripts will be released upon acceptance.
@@ 3 Bayesian Inference with Structural Missingness
Note that $X_{m}^{c}$ explicitly includes the mask $M$, ensuring that the ”observation” includes information about which variables are observed. $X_{m}$ and $X_{m}^{c}$ are still well-defined random variables to perform Bayesian inference. The formal proof that $X_{m}$ and $X_{m}^{c}$ constitute valid random variables on the appropriate disjoint union spaces is provided in Appendix C.1.
@@ 6.1 Risk Minimization and Approximation
This is proved in the Appendix Sec. C.2.
@@ 6.2 In-Model Reduction of Plug-in Imputation and MNAR Bias
This is proved in the Appendix Sec. C.3.
@@
This is proved in the Appendix Sec. C.4.
@@ 6.3 Sample Complexity Advantage of Decoupled Inference
Theorem 6.7 gives a sample-complexity advantage (ICL difficulty proxy) for learning $G$ (missing-feature posterior) and $H$ (conditional predictor) separately rather than end-to-end in finite-context / ICL. Appendix Sec. C.6 provides the supporting error decompositions and proof chain (Theorem C.9; Corollary C.10).
@@
Assume the conditional predictor regularity in Appendix Sec. C.6 so that the posterior-mismatch term is controlled by $L_{h}\,W_{1}(\cdot,\cdot)$ (Lemma C.7). Then Corollary C.10 implies that to reach total error $\epsilon$ it suffices to learn $G$ to Wasserstein accuracy $\epsilon/(2L_{h})$ and $H$ to conditional-prediction error $\epsilon/2$, capturing the required error propagation.
@@
Thus, when $L_{g}$ is large and context is limited (large $L_{\mathrm{cpl}}$), learning $G$ and $H$ separately can be more sample-efficient than learning the coupled end-to-end map; see Appendix Sec. C.6.
@@ 7 Experiments
XGBoost + MOT-MLP $0.8073\pm 0.1216$ 17.60 $132.754\pm 0.776$
PFN-NSM $0.8065\pm 0.1219$ 16.20 $0.061\pm 0.003$
XGBoost + IGRM $0.8048\pm 0.1127$ 17.90 $11.809\pm 0.156$
@@ Appendix A Related Work
A.1 Structural Missingness
While classic MCAR/MAR/ MNAR definitions (Rubin, 1976) cover probabilistic missingness, structural missingness deals with logically undefined values. Recent works have modeled this using graph-based approaches or separate categories, but often lack a unified Bayesian treatment. (Mohan et al., 2013) proposes and proved that dealing with missing data is a causal inference problem in a unified identification framework. However, due to lack of integration of methods and tools, models with simplified assumptions without considering selection bias caused by not explicitly differentiating $P(x|M=0)$ and $P(x|M=1)$ still dominate the field with plausible performance (Muzellec et al., 2020; Vo et al., 2024; Zhang et al., 2025). Though the structure is not identifiable in general without extra assumptions, PFNs can offer an “almost-free-lunch” in terms of efficient amortized inference within an assumed SCM prior: when the prior renders the relevant posteriors identifiable in-model (Remark 6.6), PFNs can approximate these posteriors and enable practical inference, rather than resolving fundamental MNAR non-identifiability without assumptions.
@@ A.1 Structural Missingness
A.2 Prior-Fitted Networks
PFNs (Müller et al., 2022, 2025) and TabPFN (Hollmann et al., 2022) demonstrate the power of learning Bayesian inference on tabular data. These methods have shown state-of-the-art performance on many small or medium-sized datasets since TabPFN-v2 (Hollmann et al., 2025). As concerns about scalability to larger datasets are raised, TabICL (Qu et al., 2025) proposes a distribution-aware feature embedding method to improve scalability, and BETA (Liu and Ye, 2025) adapts a lightweight encoder to align TabPFN to large-scale high-dimensional tabular data. As shown in the technical report for TabPFN-2.5 (Grinsztajn et al., 2025), it achieves dominant performance (87% win rate) on tabular datasets up to 100K samples and 2K features. PFN-based methods also have differentiable properties that can be leveraged to produce feature-importance measures, partial dependence (Rundel et al., 2024), and high-dimensional Bayesian optimization (Müller et al., 2023), and show promising results in settings requiring causal mechanisms, such as causal effect estimation (Balazadeh et al., 2025) and causal discovery (Robertson et al., 2025).
A.3 Generative Imputation
In tabular imputation, we aim to sample from a conditional distribution of missing entries given observed entries and a missingness mask, under heterogeneous feature types (continuous, discrete, categorical) and strong cross-feature constraints; in many practical settings this conditional mapping is close to deterministic (e.g., structurally constrained columns and rule-like relations).
@@ B.1 Missing Data Classification Experiments
Dataset selection and benchmark name.
To construct a reliable benchmark with real missingness, we filter OpenML tabular classification datasets using two simple rules: (i) the empirical missing rate is strictly positive (missing rate >0>0), and (ii) the dataset falls within the practical operating range of TabPFN (e.g., feature dimensionality and sample size constraints for feasible inference). We call the resulting benchmark the OpenML Missingness Benchmark (OMB).
We summarize the classification datasets in Tables 4 and 5, which are selected from OpenML and grouped by empirical missing rate, including dataset size (NN), inferred feature dimension ($D$), and missingness statistics (total missing entries and overall missing rate).
@@ B.2 Missing Data Imputation Experiments
XGBoost + MOT-MLP $0.8073\pm 0.1216$ 17.60 $132.754\pm 0.776$
PFN-NSM $0.8065\pm 0.1219$ 16.20 $0.061\pm 0.003$
XGBoost + IGRM $0.8048\pm 0.1127$ 17.90 $11.809\pm 0.156$
@@
TabPFN + MOT-LIN $0.8011\pm 0.1290$ 23.30 $127.294\pm 3.778$
TabPFN + MOT-MLP $0.8006\pm 0.1311$ 23.50 $128.127\pm 0.776$
TabPFN + TDM $0.8006\pm 0.1294$ 24.80 $123.860\pm 8.327$
CatBoost + Remasker $0.8005\pm 0.1298$ 27.80 $551.288\pm 9.349$
@@
TabPFN + MissForest $0.8002\pm 0.1272$ 23.20 $26.122\pm 1.681$
TabPFN + HyperImpute $0.8001\pm 0.1285$ 23.10 $170.758\pm 0.510$
@@
XGBoost + MissForest $0.7992\pm 0.1190$ 22.10 $30.749\pm 1.682$
CatBoost + IGRM $0.7986\pm 0.1155$ 25.40 $44.152\pm 0.241$
TabPFN (Raw) $0.7985\pm 0.1269$ 23.50 $0.061\pm 0.003$
TabPFN + MCFlow $0.7983\pm 0.1214$ 25.80 $28.563\pm 0.583$
XGBoost + HyperImpute $0.7982\pm 0.1281$ 23.50 $175.385\pm 0.510$
TabPFN + OTM-RR $0.7972\pm 0.1272$ 28.20 $5.241\pm 0.101$
@@
XGBoost + MCFlow $0.7935\pm 0.1187$ 21.40 $33.190\pm 0.584$
CatBoost + MCFlow $0.7901\pm 0.1161$ 25.20 $65.533\pm 0.612$
@@
Appendix C Theoretical Analysis
C.1 The Random Variables are Well-Defined
@@
C.3 Assumptions and Proof of Theorem 6.3
Assumption C.1 (Posterior approximation and predictor regularity).
@@
We provide a fully detailed proof of Theorem 6.3. Throughout this section, we fix $(X_{m}^{c},D_{m}^{c})$ and use the shorthand $\mu:=P^{\star}(\cdot\mid X_{m}^{c},D_{m}^{c})$ and $\nu:=Q_{\theta}(\cdot\mid X_{m}^{c},D_{m}^{c})$ introduced in Assumption C.1.
Lemma C.2 (Tower property representation).
@@
Lemma C.3 (Error decomposition).
@@
By Lemma C.2, $T_{g}^{\star}=\mathbb{E}_{\mu}[h^{\star}(X_{m})]$. Therefore,
@@
Lemma C.4 (Lipschitz test functions and Wasserstein-1).
@@
Assumption C.1(A0) ensures $W_{1}(\mu,\nu)$ is well-defined. By Kantorovich–Rubinstein duality,
@@
The second term is bounded by Assumption C.1(A3), yielding $\left|\mathbb{E}_{\nu}[h^{\star}-\hat{h}_{\phi}]\right|\leq\varepsilon_{\text{pred}}$. For the first term, Lemma C.4 gives $\left|\mathbb{E}_{\mu}[h^{\star}]-\mathbb{E}_{\nu}[h^{\star}]\right|\leq L\,W_{1}(\mu,\nu)$, and Assumption C.1(A2) implies $W_{1}(\mu,\nu)\leq\varepsilon_{\text{post}}$. Combining yields
@@
C.6 Posterior-Predictive Approximation via Decoupled Heads
@@
Lemma C.5 (Posterior-predictive marginalization).
@@
Lemma C.6 (Decoupled approximation error decomposition).
@@
Lemma C.7 (Bounding posterior mismatch by $W_{1}$).
Assume that for fixed $(x_{m}^{c},D_{m}^{c})$ and for any $\varphi$ with $\|\varphi\|_{\infty}\leq 1$, the map $x_{m}\mapsto\mathbb{E}_{y\sim H(\cdot\mid x_{m},x_{m}^{c},D_{m}^{c})}[\varphi(y)]$ is $L_{h}$-Lipschitz w.r.t. $\|\cdot\|_{2}$. Then the first term in Lemma C.6 is bounded by
@@
Lemma C.8 (From conditional KL to bounded test functions (Pinsker)).
@@
Proposition C.9 (Why decoupling helps approximate Bayes posterior predictive with limited context).
Under the identifiability condition in Remark 6.6 (so that $P^{\star}(x_{m}\mid x_{m}^{c},D_{m}^{c})$ is well-defined within the model class), Lemmas C.6C.7 show that approximating the Bayes posterior predictive reduces to separately learning (i) the posterior over missing features ($G$) and (ii) the conditional predictor ($H$). Moreover, Lemma C.8 shows how PFN-style training that minimizes conditional cross-entropy (equivalently, conditional KL in the realizable case) controls bounded test-function error for the marginalized posterior predictive without explicitly computing the integral in Lemma C.5. Combined with Theorem 6.7, this supports the claim that, with limited context (test-time training / in-context learning), a decoupled multi-task design can be more sample-efficient and thereby more favorable for approximating the Bayes-optimal posterior predictive than directly learning the coupled end-to-end map.
@@
Under Remark 6.6, the Bayes posterior predictive is well-defined in-model. Lemmas C.6C.7 decompose the approximation error into (i) a posterior-mismatch term (bounded by a $W_{1}$ distance for $G$) and (ii) a conditional-prediction term for $H$. Hence, approximating the posterior predictive reduces to learning $G$ and $H$ to control these two terms. Lemma C.8 provides a direct route for the PFN head: small conditional KL to the true posterior predictive implies small bounded test-function error for the marginalized predictor. Theorem 6.7 adds a complementary complexity argument: when the structural component is highly nonlinear (large effective $L_{g}$) and context is limited (large coupling difficulty), learning the decoupled components can be more sample-efficient than learning the coupled end-to-end map in the ICL regime. ∎
Corollary C.10 (Explicit link to complexity via a two-term target).
Under the assumptions of Lemmas C.6C.7, for any $\varphi$ with $\|\varphi\|_{\infty}\leq 1$,
@@
Apply Lemma C.6 and then upper bound the first (posterior approximation) term by Lemma C.7. The remaining term is exactly $\epsilon_{H}(x_{m}^{c},D_{m}^{c})$ by definition. ∎
C.7 Proof of Theorem 6.7
@@
By Corollary C.10, to ensure total test-function error at most $\epsilon$ it suffices to satisfy
@@
Figure 4: Imputation benchmark under MCAR missingness: out-of-sample MAE over 15 datasets, plus AVG and RANK (mean $\pm$ std over masks).
Appendix D Additional Experiments
We report additional imputation experiments under the MCAR missingness mechanism.
Figure 4 summarizes out-of-sample MAE across the same 15 imputation benchmark datasets, together with the overall average (AVG) and the average rank (RANK).
We additionally compare (i) our method against boosting baselines, (ii) TabPFN/PFN-family variants, and (iii) top split baselines, in that order.
D.1 Regression results under MCAR missingness
Across the 15 datasets in Figure 4, PFN-Flow is consistently competitive under MCAR missingness. In particular, PFN-Flow tends to achieve either the best or near-best MAE on a majority of datasets, and remains in the top tier in terms of AVG and RANK, indicating both strong average performance and stable behavior across datasets and masks. We also observe that some baselines can be highly dataset-dependent (large performance variance across datasets), whereas PFN-Flow exhibits comparatively uniform improvements.
D.2 Boosting baselines
Figure 5 compares PFN-Flow with three standard boosting baselines (XGBoost/CatBoost/LightGBM, trained on the same MCAR-imputed data). Each scatter contains two panels corresponding to missing rate $\leq 10\%$ and $10\%$; for each dataset, the x-axis is PFN-Flow AUC and the y-axis is the baseline AUC. Points below the diagonal therefore indicate that PFN-Flow outperforms the baseline on that dataset/group, while points above the diagonal indicate the opposite. Overall, most points lie on or below the diagonal, suggesting that PFN-Flow remains competitive against strong non-neural baselines under MCAR, with only a small number of datasets where boosting can be slightly better.
D.3 TabPFN variants and PFN-family baselines
Figure 6 further compares PFN-Flow to TabPFN (Raw) and a PFN-family baseline (PFN-NSM). We observe that the points are highly concentrated near the diagonal in both missingness regimes, indicating that PFN-Flow is broadly comparable to TabPFN/PFN-family methods on the same datasets under MCAR. Notably, in the higher-missingness group ($10\%$), the scatter remains close to parity with only a few outliers, suggesting that PFN-Flow does not rely on a narrow set of favorable datasets and maintains stable performance as missingness increases.
D.4 Top split baselines
Figure 7 reports the strongest split baselines selected from our splitting-based design space (Top-1 to Top-4), each corresponding to a pipeline of a predictive model (here, CatBoost) and a specific imputation module. Within each subfigure, the two internal panels again correspond to missing rate $\leq 10\%$ and $10\%$. For each dataset, the x-axis is PFN-Flow AUC and the y-axis is the baseline pipeline AUC; points below the diagonal indicate that PFN-Flow improves upon that strong split baseline. Across all four top pipelines, the majority of points remain on or below the diagonal, showing that PFN-Flow is competitive even against the best-performing split configurations. At the same time, the few points above the diagonal highlight that no single method universally dominates every dataset, motivating the need for robust imputation that generalizes across heterogeneous tabular domains.
(a) XGBoost (Raw) vs PFN-Flow (AUC).
(b) CatBoost (Raw) vs PFN-Flow (AUC).
(c) LightGBM (Raw) vs PFN-Flow (AUC).
Figure 5: Boosting baseline comparisons against PFN-Flow under MCAR. Each point corresponds to one dataset in one missingness group; the diagonal indicates parity.
(a) TabPFN (Raw) vs PFN-Flow (AUC).
(b) PFN-NSM vs PFN-Flow (AUC).
Figure 6: TabPFN/PFN-family comparisons against PFN-Flow under MCAR. Each point corresponds to one dataset in one missingness group; the diagonal indicates parity.
(a) Top-1 split: CatBoost + GRAPE.
(b) Top-2 split: CatBoost + MOT-MLP.
(c) Top-3 split: CatBoost + Remasker.
(d) Top-4 split: CatBoost + OTM-SK.
Figure 7: Top split comparisons.